# Finite phase coherence time of a quantum field created by an ideal Bose gas

### v5

@@ Abstract
To study the relaxation to thermal equilibrium (in the lower panel of Fig. 2), the conditional probability condition in Eq. (4) is imposed additionally to the constraint of particle number conservation in Eq. (5) and the condition of constant temperature in the Monte Carlo sampling algorithm for the time scales $t_{1}$ and $t_{2}$ in Eq. (3). From these two constraints and the sampling of complex time, one may calculate and analyze the absolute value of complex time, which corresponds to the single particle entropy (finite quantum coupling to the atomic vapor), related to both the coherence time scale $t_{2}$ and the oscillation time scale $t_{1}$, as shown in Fig. 2 for conditional probabilities. From Figs. 2, 3 one can extract that the absolute value of the time measure $t$ is distributed around a mean value of approximately $(405\pm 15)~{}\mu s$ at a given typical temperature of $T=5.0$ nK. The distribution of the absolute value of time $t$ has a Gaussian shape and tends to be always distributed around finite quantitative values as compared to no conditional probability (not shown). This smoothing is due to the condition that the system has to satisfy the ergodicity assumption of the Boltzmann equilibrium in the present model, where not only the condition of constant temperature and the conservation of particle number, but also the condition of a stable Boltzmann equilibrium is imposed onto the applied sampling method or algorithm, respectively.
Calculating the entropy in units of the Boltzmann constant at constant temperature highlights Gaussian fluctuations (below as well as close above the critical temperature for Bose-Einstein condensation) in Fig. 4. Starting from temperatures not too far above the critical temperature, where the fugacity spectrum obeys a gap between the (symmetric) equilibrium state and the excited states (which are symmetry breaking), from the scaling of the entropy with temperature (not shown graphically), it can be understood that the process of spontaneous symmetry breaking is due to these non-zero fluctuations of the quantum entropy (particle coherence among different single-particle quantum states), occurring as a purely quantum mechanical effect which decays on the time scale $t_{2}<\hbar\beta$ and does not require interactions to the lowest order perturbation theory of the present setup. Thus, the definition of a well-defined gauge for the quantum field ($\psi=0$) and the Boltzmann equilibrium is not only a necessary, but also a sufficient condition for breaking phase gauge symmetry of the quantum field in the given parameter regime. In the present study, this is not the case for a uniform distribution, since as indicated numerically, the finite coherence time can tend to zero in the case of a uniform distribution (not shown explicitly). However, this does not entirely prove that the Boltzmann equilibrium itself is a necessary condition for Bose-Einstein condensation, but only that defining a gauge is a necessary and sufficient condition for Bose-Einstein condensation due to finite coherent times - if conditional probabilities are assumed to follow the thermal Boltzmann equilibrium. Finally, the fact that the entropy is maximized and finite at thermal equilibrium indicates the non-classical correlations due to quantum coherence on the (scaled) time scale $t_{2}$ also at the thermal Boltzmann equilibrium.
Figure 4: (color online) Figure shows $5*10^{5}$ realizations of the maximum entropy (absolute value of the distribution of complex time $t$ in units of the Boltzmann constant $k_{B}$, compare Eq. (4)) for a quantum particle in a Bose-Einstein condensate from Monte Carlo sampling using the ergodicity assumption. A Gaussian distribution is observed for non-isotropic trap geometries $\omega_{x},\omega_{y}=2\pi\times 42.0$ Hz and $\omega_{z}=120.0$ Hz and temperature $T=5.0$ nK, indicating the randomness of the quantum field and its fluctuations at thermal equilibrium.
@@
In the present theory, it is illustrated that the constraint of particle number conservation naturally leads to the composition of real-valued and purely imaginary valued time to an extended complex-valued number, which defines time in terms of a two-dimensional complex variable to fully capture the system in its full complexity - interacting particles below the critical temperature. It is the spectrum of this complex time with the corresponding fugacity spectrum that naturally defines the direction of a composite (complex-valued) time, that is two simultaneous processes described by $t_{1}$ which describes coherent oscillations of the quantum particle’s wave function and $t_{2}$ which defines the (de-)coherence time of a single quantum particle in the presence of the $N-1$ other particles, in the form of a cone propagating in the positive imaginary direction of complex space on average. In that representation, i. e. after coupling a single particle to the rest of the gas within the time scale $t_{2}\leq\hbar\beta$, complex time doesn’t obey time-reversal symmetry anymore, indicating the spontaneous breaking of the gauge symmetry for the quantum field, defined in terms of a variable for the chemical potential which is proportional to complex time. In the sequel of a quantitative analysis of complex time, it turns out that the real part of time is typically distributed around zero (that is almost all particles share the same coherent phase), whereas the imaginary part of complex time mostly entails the system’s relaxation properties (finite coherence times, because of spontaneous fluctuations and decoherence). In the limit of vanishing temperatures, the process of spontaneous symmetry breaking (characterized by the well-defined relative phase relation between condensate and non-condensate particles) is directly and reproducibly measurable as condensate formation, respectively, as reported for instance in Ref. [5]. Measuring the relative phase between possible condensate and non-condensate states of a quantum particle in the Bose-Einstein condensate, especially in an intermediate temperature range below zero and the critical temperature, Bose-Einstein condensation in nearly-ideal Bose gases is direct evidence for spontaneous symmetry breaking from pure quantum fluctuations, as discussed in the sequel of the present theory.
It is the absolute value of complex time, which defines the entropy of the quantum particle in the Bose-Einstein condensate in units of the Boltzmann constant $k_{B}$. The finite value as well as the width of the entropy distribution indicates the coupling of several quantum states which do not belong to an exact Boltzmann equilibrium state as originally predicted in the framework of classical (equilibrium) many-particle mechanics, but are modified by the uncertainty and fluctuations of the particle’s entropy in the atomic cloud at finite temperature. The real-valued part of time $t_{1}$ is directly related to the fugacity ring spectrum in the complex plane, whereas the coherence time $t_{2}$ describes the coupling of quantum states in the symmetry-breaking part of the fugacity spectrum. At the Boltzmann equilibrium, the distribution of entropy values ideally occurs in connection to quantum states in the neighborhood of ${\rm Re}(z)=1$ and ${\rm Im}(z)=0$ (fugacity equals one). Thus, the finite value of the entropy and its uncertainty indicates and quantifies the coupling of symmetric and asymmetric parts of the fugacity spectrum, which thus appear only structurally unconnected above the critical temperature. Finally, the different shape of the entropy distribution from a flat distribution with a full range from ${\rm Re}(z)=0$ to ${\rm Re}(z)=1$ to a Gaussian distribution for conditional Boltzmann probabilities indicates that ergodicity leads to finite coherence times in the Boltzmann limit of maximum entropy, ensuring the zero gauge of the total spatially integrated quantum field at spatially uniform relative phase distributions of the single-particle wave functions for the quantum field to be a sufficient condition for Bose-Einstein condensation.
@@
In conclusion, using a representation of complex time, the coherence time of a quantum particle in a Bose-Einstein condensate is on the order of a few hundred microseconds in the typical parameter regime of Bose gases at thermal equilibrium in harmonic traps below the critical temperature. In the presented quantum field theory, the definition of a complex-valued time arises naturally from the constraint of particle number conservation, mathematically defined by complex poles of an intrinsic equation for the fugacity (and the chemical potential) of the Bose-Einstein condensate. Within the presented model, gauging the quantum field in the Boltzmann equilibrium is a necessary and sufficient condition for spontaneous symmetry breaking for a nearly ideal Bose gas. Moreover, it is possible to argue and numerically verify that the process of spontaneous symmetry breaking is induced by time and entropy fluctuations enhanced by the reduction of the atomic density below the critical density for Bose-Einstein condensation, which couples (unconnected) symmetric and asymmetric parts of the fugacity spectrum. In the classical limit, entropy fluctuations tend to be distributed around a ring in the complex spectrum of the fugacity, recovering the time-reversal symmetry of the quantum field.

### v6

@@ Abstract
To study the relaxation to thermal equilibrium (in the lower panel of Fig. 2), the conditional probability condition in Eq. (4) is imposed additionally to the constraint of particle number conservation in Eq. (5) and the condition of constant temperature in the Monte Carlo sampling algorithm for the phases $\Gamma_{1}t$ and $\Gamma_{2}t$ in Eq. (3). From these two constraints and the sampling of complex time, one may calculate and analyze the absolute value of complex time, which corresponds to the single-particle entropy (finite quantum coupling to the atomic vapor), related to both the coherence time scale $\Gamma_{2}t$ and the oscillation time scale $\Gamma_{1}t$, as shown in Fig. 2 for conditional probabilities. From Figs. 2, 3 one can extract that the absolute value of the time measure $t$ is distributed around a mean value of approximately $(405\pm 15)~{}\mu s$ at a given typical temperature of $T=5.0$ nK. The distribution of the absolute value of time $t$ has a Gaussian shape and tends to be always distributed around finite quantitative values as compared to no conditional probability (not shown). This smoothing is due to the condition that the system has to satisfy the ergodicity assumption of the Boltzmann equilibrium in the present model, where not only the condition of constant temperature and the conservation of particle number, but also the condition of a stable Boltzmann equilibrium is imposed onto the applied sampling method or algorithm, respectively.
@@
In the present theory, it is illustrated that the constraint of particle number conservation naturally leads to the composition of real-valued and purely imaginary valued time to an extended complex-valued number, which defines time in terms of a complex variable to fully capture the system in its full complexity - interacting particles below the critical temperature. It is the spectrum of this complex time with the corresponding fugacity spectrum that naturally defines the direction of a composite (complex-valued) time, that is two simultaneous processes described by $\Gamma_{1}t$ which describes coherent oscillations of the quantum particle’s wave function and $\Gamma_{2}t$ which defines the (de-)coherence time of a single quantum particle in the presence of the $N-1$ other particles, in the form of a very tight cone propagating in the positive imaginary direction of complex space on average. In that representation, i. e. after coupling a single particle to the rest of the gas within the time scale $\Gamma_{2}t\leq\hbar\beta$, complex time doesn’t obey time-reversal symmetry anymore. In the sequel of a quantitative analysis of complex time, it turns out that the real part of weighted complex time is typically distributed around zero (that is almost all particles share the same coherent phase), whereas the imaginary part of weighted complex time mostly entails the system’s relaxation properties (finite coherence times, because of spontaneous fluctuations and decoherence). In the limit of vanishing temperatures, the process of spontaneous symmetry breaking (characterized by the well-defined relative phase relation between condensate and non-condensate particles) is directly and reproducibly measurable as condensate formation, respectively, as reported for instance in Ref. [5]. Measuring the relative phase between possible condensate and non-condensate states of a quantum particle in the Bose-Einstein condensate, especially in an intermediate temperature range below zero and the critical temperature, Bose-Einstein condensation in nearly-ideal Bose gases is direct evidence for spontaneous symmetry breaking from quantum fluctuations below the critical temperature.
@@
In conclusion, using a representation of complex time, the (average) coherence time of a quantum field created by a Bose-Einstein condensate is on the order of a few hundred microseconds in the typical parameter regime of Bose gases at thermal equilibrium in harmonic traps below the critical temperature. The phase coherence of the quantum field calculated within the present theoretical model can be understood as the spatially averaged atomic field that shows temporal phase correlations. Such phase coherence may e.g. be experimentally obtained from the (pulsed) measurement of reflected laser light of the atomic sample. In the presented quantum field theory, the coherence time defines a unit scale for the forward propagation of time from destructive partial wave interference and a decay of the average quantum field, whereas the direction of time is due to the equlibration of the Bose gas to the Boltzmann equilibrium.
Within the presented model, defining a fixed gauge by the Boltzmann equilibrium is a necessary and sufficient condition for spontaneous symmetry breaking in the Bose gas - while each realization of a Boltzmann equilibrium is stable on the coherence time scale of a few hundred microseconds as presented. From the spectrum of the fugacity, it is possible to understand and numerically verify that the process of spontaneous symmetry breaking is induced by energy fluctuations that are induced by the reduction of the atomic density below the critical density for Bose-Einstein condensation, which couples (unconnected) symmetric and asymmetric parts of the fugacity spectrum. In the classical limit, entropy fluctuations tend to be distributed around a ring in the complex spectrum of the fugacity, approaching configurations of time-reversal symmetry for the quantum wave field.
@@
Acknowledgements.
The author acknowledges the financial support from IU Internationale Hochschule for the (freelancer) lecturer position at the university, which has in particular enabled the formulation and editing of the present theory on the finite phase coherence time of a quantum field created by an ideal Bose gas.